OpenStudy (anonymous):

Solve for x: -3|x + 7| = -12

OpenStudy (anonymous):

Those lines just mean that the x+7 is positive so just think of it as parentheses. Distribute the -3 in (x+7).

OpenStudy (anonymous):

You should get -3x-21.

OpenStudy (anonymous):

-3x-21=-12

OpenStudy (anonymous):

Now try to solve this.

OpenStudy (anonymous):

It should be easier.

OpenStudy (anonymous):

x-18=-12

OpenStudy (anonymous):

right???

OpenStudy (anonymous):

Ya.

OpenStudy (anonymous):

x=6

OpenStudy (anonymous):

Then the answer would be x=6. But you aren't done there. When I told you that the absolute value is like parentheses, I was trying to not confuse you by saying that now you have to put a negative in front of the parentheses.

OpenStudy (anonymous):

Absolute values will always be positive and negative when it comes to equations.

OpenStudy (freckles):

you should get two answers first divide both sides by -3 |x+7|=4 you have to solve both equations below: x+7=4 or x+7=-4

OpenStudy (anonymous):

So now it should look like this(-3){-(x+7)}=-12

OpenStudy (anonymous):

then both by -11

OpenStudy (anonymous):

-3(-x-7)=-12 3x+21=-12 3x=-21-12 3x=-33 x=-11

OpenStudy (anonymous):

Yes that's the second answer.

OpenStudy (freckles):

x=6 gives a false equation -3|6+7|=-12 doesn't hold -3|13|=-12 -3(13)=-12 -39=-12 see this isn't the same thing on both sides so x=6 isn't a correct solution

OpenStudy (anonymous):

thxs i got my answers you guys are the best

OpenStudy (anonymous):

I don't think it's suppose to equal back to the problem @freckles. Look at this link. http://www.purplemath.com/modules/solveabs.htm

OpenStudy (anonymous):

And you're welcome!

OpenStudy (freckles):

yes it is in order for x=c to be a solution it must satisfy the equation

OpenStudy (freckles):

the first equation was just solved incorrectly

OpenStudy (anonymous):

When you plug in the 1st x of the equation from the picture. It doesn't equal 7. The second x does though.

OpenStudy (freckles):

-3(x+7)=-12 -3x-21=-12 add 21 on both sides -3x=-12+21 -3x=21-12 -3x=9 divide by -3 on both sides x=-3 as you see pluggin x=-3 into original this does satisfy initial equation -3|x+7|=-12 replace x with -3 -3|-3+7|=-12 -3|4|=-12 -3(4)=-12 -12=-12 this is a true equation so x=-3 is a solution just like x=-11 is also a solution since -3|x+7|=-12 pluggin in -11 for x -3|-11+7|=-12 -3|-4|=-12 -3(4)=-12 -12=-12 same thing on both sides the equation is true x=-11 is also a solution so the solution set is {-3,-11}

OpenStudy (freckles):

@200205650 you have to plug into the original equation

OpenStudy (anonymous):

Okay then never mind. You are right.